40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 14 20:42:04 2021, gautam, Update, LSC, Locking 
    Reply  Thu Jul 15 14:25:48 2021, Paco, Update, LSC, CM board high_BW_TFs.pdf
Message ID: 16248     Entry time: Thu Jul 15 14:25:48 2021     In reply to: 16247
Author: Paco 
Type: Update 
Category: LSC 
Subject: CM board 

[gautam, paco]

We tested the CM board by implementing the high bandwidth IR lock (single arm). In preparation for this test we temporarily connected the POY11_Q_MON output to the CM board IN1 input and checked the YARM POY transfer function by running the AA_YARM_TEMPLATE under users/Templates/LSC/LSC_loops/YARM_POY/. We made sure the YARM dither optimized TRY so as to maximize the optical gain stage. Then we proceeded as follows:

  • From the LSC --> CM Servo screen, we controlled the REFL 1 Gain (dB) slider (nominal +25) and MC Servo IN2 Gain (dB) slider (nominal -32 dB) to transfer the low bandwidth (digital) control to the high bandwidth (analog) control of the YARM.
  • During this game, we monitored the C1:LSC-POY11_I_ERR_DQ & C1:LSC-CM_SLOW_OUT_DQ error signal channels for saturation, oscillations, or stability.
  • Once a set of gains was successful in maintaining a stable lock, we measured the OLTF using SR 785 to track the UGF as we mix the two paths.
  • Once the gains have increased, a boost and super-boost stages may be enabled as well.

Ultimately, our ability to progressively increase the control bandwidth of the YARM is a proxy that the CM board is working properly. Attachment 1 shows the OLTF progression as we increased the loop's UGF. Note how as we approached the maximum measured UGF of ~ 22 kHz, our phase margin decreased signifying poor stability.


At the end of this measurement, at about ~ 15:45 I restored the CM board IN1 input and disconnected the POY11_Q_MON

gautam: the conclusion here is that the CM board seems to work as advertised, and it's not solely responsible for not being able to achieve the IR handoff. 

Attachment 1: high_BW_TFs.pdf  43 kB  Uploaded Thu Jul 15 16:45:02 2021  | Show | Show all
ELOG V3.1.3-